# Not-Just-Scaling Laws: Towards a Better Understanding of the Downstream Impact of Language Model Design Decisions

### v1

@@ Abstract
Improvements in language model capabilities are often attributed to increasing model size or training data, but in some cases smaller models trained on curated data or with different architectural decisions can outperform larger ones trained on more tokens. What accounts for this? To quantify the impact of these design choices, we meta-analyze 92 open-source pretrained models across a wide array of scales, including state-of-the-art open-weights models as well as less performant models and those with less conventional design decisions. We find that by incorporating features besides model size and number of training tokens, we can achieve a relative 3-28% increase in ability to predict downstream performance compared with using scale alone. Analysis of model design decisions reveal insights into data composition, such as the trade-off between language and code tasks at 15-25% code, as well as the better performance of some architectural decisions such as choosing rotary over learned embeddings. Broadly, our framework lays a foundation for more systematic investigation of how model development choices shape final capabilities. [^1] Code and data are available at https://github.com/nightingal3/llm-pretraining-behaviours for the community to build upon.
@@
Emmy Liu1, Amanda Bertsch1, Lintang Sutawika1, Lindia Tjuatja1, Patrick Fernandes1,2,3 Lara Marinov1, Michael Chen1, Shreya Singhal1, Carolin Lawrence4, Aditi Raghunathan1, Kiril Gashteovski4, 5, Graham Neubig1 1 Carnegie Mellon University, Language Technologies Institute 2Instituto Superior Técnico (Lisbon ELLIS Unit), 3 Instituto de Telecomunicações 4 NEC Laboratories Europe, Germany 5 Center for Advanced Interdisciplinary Research, Ss. Cyril and Methodius Uni. of Skopje, Germany Correspondence: emmy@cmu.edu
@@ 3 Building a Database of Publicly-Available Language Models
To approach our research question, we built a comprehensive database of publicly available language models. Our database encompasses models spanning a wide range of sizes, from 11M to 110B parameters, [^3] Including embedding parameters. and includes only distinct pretrained base models with decoder-only architectures. [^4] By distinct, we mean that each combination of training data and model architecture decisions should be unique. Variants of the same model trained on a deduplicated dataset are counted as separate, but not variants trained on the same data but with different curricula/initializations. In this section, we describe the criteria used for model inclusion, how we featurized the models, as well as the evaluation suite we used.
@@ 3.2.2 Exploring Data Composition via Generation
Although many models do document some details of their data composition, relatively few release their full pretraining corpus. Further, even when data composition statistics are provided, they are often presented at a high level of granularity. As a result, many models in our study have missing values for data composition.
To address this, we further propose an approach of estimating the data a model was trained on by generating from the model with a context containing just the beginning-of-sequence (BOS) token for that model (or end-of-sequence if that model lacked a BOS token). We use temperature-based sampling with $T=1$, as this should in principle recover $P_{raw}(x_{t}\mid x_{<t})\approx P(x_{t}\mid x_{<t})$ in the limit of sampling infinitely from an LM that captures its distribution perfectly.
We call this free-generation. However, this has important caveats: we cannot actually sample infinitely, and a model does not reflect its pretraining data perfectly. Therefore, we use this as a “fingerprint” of the model’s training, but do not claim that this reflects a model’s pretraining dataset exactly. In practice, we sample between 5-10k free-generations from each model (5k for larger models). In addition to categorizing them (see Appendix E) by the domains in Figure 2, we also extract lower-level features:
Low-level data features:
We aggregate per-generation statistics that reflect data quality and linguistic structure. Examples include the words per sentence, the depth of the constituency tree for natural language text, and dependency length.
@@ 3.3 Evaluation Datasets and Metrics
Table 1: Overview of LM evaluation datasets with approximate sample counts, citations, and evaluation metrics. Datasets ANLI, XNLI, LogiQA2, and MathQA use Brier Score, while the others use Accuracy.
To assess how design choices affect reasoning capabilities, we evaluated model performance on a curated suite of datasets derived from the first version of the Open LLM leaderboard Myrzakhan et al. (2024), which was designed to capture diverse facets of reasoning (Table 1). [^5] Arithmetic and Minerva math (Brown et al. (2020); Hendrycks et al. (2021)) tasks were also initially included in this leaderboard, but we excluded them as we focused solely on base (not instruction tuned) models, and very few were able to achieve non-zero scores. We collect results for some models directly from the leaderboard, and for models not on the leaderboard we use the Eleuther LM eval harness (Gao et al., 2023) to conduct evaluations with exactly the same setting. In addition, if there were multiple versions of a task or sub-tasks, we evaluated all of them and averaged them to get the overall task score. For the full list of evaluation datasets and settings, see Appendix C.
@@
For tasks where smaller models struggle to achieve non-zero accuracy, we follow Schaeffer et al. (2023) in using multiclass brier score as an alternate continuous metric for multiple-choice tasks Brier (1950). For a task with $K$ classes, let $p_{ik}$ be the predicted probability for class $k$ on sample $i$. Then $s_{T,BS}=\frac{1}{|T|}\sum_{i=1}^{|T|}\sum_{j=1}^{K}(p_{ik}-\mathds{1}{\{y_{i}=k\}})$. [^6] Note that lower is better for brier score. Multiclass brier score ranges between 0-2.
@@ 4 Predictive Modeling
Next, given our database we fit a regressor to try to predict performance. In traditional scaling laws, regressors are fit based on power laws. However, we are now dealing with a larger number of features, some of which may not be captured well by simple parametric forms. Hence, we follow previous work on performance prediction (Xia et al., 2020; Ye et al., 2021) utilizing tree-based regressors based on XGBoost Chen and Guestrin (2016). [^7] We also performed preliminary experiments with LightGBM Ke et al. (2017) but it yielded very similar results in both prediction accuracy and feature importance.
For each evaluation benchmark, we train a model to predict the performance metric on that task based on architectural features $\mathcal{A}$ and data features $\mathcal{D}$. For each task setting, we perform 3-fold cross-validation due to the relatively small number of models, with a nested inner cross-validation over the training set in each fold. The inner cross-validation conducted grid search over a small set of hyperparameters, allowing the model to slightly vary per task. See Appendix G for more details.
@@
To evaluate the predictors, we use Mean Absolute Error averaged across all models and folds. In other words, for a task with $N$ models evaluated, $\text{MAE}_{T}=\frac{1}{|T|}\sum_{i=1}^{N}|s_{T}(M_{i})-\widehat{s_{T}}(M_{i})|$. We compare the scaling-laws predictor as well as the all-features predictor against each other, but also against the median baseline, which simply predicts the median score of the models in the training set for each model in the test set of that fold.
@@
As the full set of features is very large, we sequentially selected features from the full set greedily based on which reduced MAE the most, averaged across 5 random seeds. Features were added until no reduction of at least $1\times 10^{-4}$ was observed. We started using only the two scaling laws features, and refer to this as the scaling-laws model, though it does not have the form of a traditional power law.888 Typically, scaling laws are used to extrapolate the performance of larger models. Because we use a decision-tree based predictor, our approach is less likely to extrapolate, a trade-off we opted to take to incorporate the array of scale-independent features we have, not all of which are numeric. Therefore, we moreso focus on interpolating performance within the size boundaries that we have (roughly 10M-100B parameters, and 50B-3T tokens). Examining results with a variety of other prediction methods is an interesting direction for future work. By then incorporating additional architectural or data features, we can then directly quantify the incremental predictive power afforded by these extra features. We refer to the model with the set of features as the all-features model. In all cases, we ran models with the same hyperparameter grid and the same random seeds and splits.
@@
Benchmark Setting Baseline MAE Scaling Laws MAE All Features MAE p-val (corrected)
@@
Arc Challenge 25-shot 13.23% 4.36% $\pm$ 0.12% 3.67% $\pm$ 0.09% $4.89\times 10^{-19}$
GSM8k 5-shot 15.65% 6.04% $\pm$ 0.21% 5.10% $\pm$ 0.23% $6.49\times 10^{-14}$
Hellaswag 10-shot 12.26% 3.93% $\pm$ 0.13% 3.18% $\pm$ 0.09% $6.66\times 10^{-20}$
Humaneval 0-shot 11.79% 8.08% $\pm$ 0.22% 6.93% $\pm$ 0.22% $1.46\times 10^{-12}$
Lambada 0-shot 16.89% 9.51% $\pm$ 0.33% 6.85% $\pm$ 0.25% $2.87\times 10^{-22}$
MMLU (0-shot) 0-shot 11.98% 4.76% $\pm$ 0.20% 4.10% $\pm$ 0.17% $6.02\times 10^{-13}$
MMLU (5-shot) 5-shot 12.25% 3.97% $\pm$ 0.18% 3.54% $\pm$ 0.14% $2.09\times 10^{-10}$
TruthfulQA 0-shot 3.72% 2.75% $\pm$ 0.08% 2.29% $\pm$ 0.06% $1.27\times 10^{-17}$
Winogrande 5-shot 10.14% 3.39% $\pm$ 0.08% 3.09% $\pm$ 0.07% $6.02\times 10^{-13}$
@@
XNLI 0-shot 7.22 6.68 $\pm$ 0.11% 6.30$\pm$ 0.11% $3.16\times 10^{-9}$
ANLI 0-shot 5.90 6.74 $\pm$ 0.19% 6.53$\pm$ 0.21% $3.84\times 10^{-4}$
MathQA 0-shot 7.57 2.83 $\pm$ 0.06% 2.75$\pm$ 0.07% $1.63\times 10^{-4}$
LogiQA2 0-shot 12.74 4.74 $\pm$ 0.12% 4.60$\pm$ 0.15% $1.37\times 10^{-2}$
Table 2: Comparison of MAE values (mean $\pm$ 95% CI) for Scaling Laws and All Features predictors alongside Baseline MAE. Lower MAE is bolded; * indicates significance ($p<0.05$). Brier score values are multiplied by 100 to be on a similar scale to accuracy.
@@ 5.2 What Features Does Task Performance Depend On?
To better understand the factors that influence task performance, we examine Shapley (1953) (SHAP) values of the predictor, which provide a local view of how individual feature values influence predictions. The results and feature descriptions for Arc Challenge, HumanEval, Winogrande, and TruthfulQA are shown in Figure 4, and results for remaining benchmarks are shown in Appendix I.
A little code goes a long way, but too much is harmful to NLI. One of the most important non-scaling features is the percentage of code data in pretraining. Higher code composition results in positive Shapley values (i.e. higher predicted performance) for Humaneval, but it negatively affects Arc Challenge, Hellaswag, Winogrande, and Lambada. In the scatterplots of Figure 5, we compare the code percentage against SHAP impact for both small and large models. We see that models trained with more than roughly 20–25% code are predicted to have large gains on tasks like Humaneval, but start to incur penalties on standard natural language benchmarks. By contrast, a moderate code proportion in the 15–25% range appears to balance these competing demands, yielding a more neutral or slightly positive effect overall.
@@
Other domains of data can occasionally yield task-specific effects. While the percentage of code in pretraining is consistently selected as an impactful feature, with clear trade-offs, this other pretraining domains are selected less frequently. From the fine-grained features from free generations, we also observed that many recent models (particularly those trained on synthetic data such as the Phi (Gunasekar et al., 2023) and SmolLM (Allal et al., 2024)) generate a relatively large number of question words, indicating extensive training on data related to question answering. A higher percentage of reference-like or question-loaded generations resulted in better model accuracy on some tasks such as Arc Challenge and Winogrande. Additionally, models that generate more web-like data tend to do worse on TruthfulQA (Figure 4).
@@ 6.1 Empirical Data Composition Results
Prior work has examined the effects of including code during pretraining Ma et al. (2023); Aryabumi et al. (2024) and ablating domains such as C4 or books from The Pile Longpre et al. (2024). Data filtering has also been shown to improve performance beyond scaling alone by removing low-quality data Sorscher et al. (2023); Goyal et al. (2024). Our results align with prior findings, indicating that code can enhance natural language reasoning at moderate proportions but degrades performance at higher percentages. We estimate an optimal code ratio of $15\text{-}25\%$, refining prior work suggesting $25\%$ Aryabumi et al. (2024), though intermediate ranges were not tested in their study. Our approach—first pooling insights from existing models—complements empirical ablations by identifying useful axes of variation for further testing.
@@ 6.2 Observational and Task-Specific Scaling Law Fitting
Previous works have examined task-specific scaling laws. In machine translation, parameter allocation between encoder and decoder affects outcomes, and incorporating machine-translated data can be detrimental Ghorbani et al. (2021). Multilingual studies reveal that language similarity doesn’t impact scaling trends; however, multitasking offers greater benefits when English is the target language Fernandes et al. (2023). Scaling laws for downstream tasks and transfer learning have been proposed, emphasizing that alignment between pretraining data and downstream tasks is crucial for performance prediction Hernandez et al. (2021); Isik et al. (2024). Data repetition has been considered, especially in data-constrained domains Muennighoff et al. (2024), with extensions to multiple data domains Goyal et al. (2024). Alternative scaling formulations address factors like sparsity Frantar et al. (2023), precision Kumar et al. (2024), and inference costs Hoffmann et al. (2022b). In contrast, some studies find stable performance across various batch sizes and learning rates DeepSeek-AI et al. (2024a).

### v2

@@ Abstract
Improvements in language model capabilities are often attributed to increasing model size or training data, but in some cases smaller models trained on curated data or with different architectural decisions can outperform larger ones trained on more tokens. What accounts for this? To quantify the impact of these design choices, we meta-analyze 92 open-source pretrained models across a wide array of scales, including state-of-the-art open-weights models as well as less performant models and those with less conventional design decisions. We find that by incorporating features besides model size and number of training tokens, we can achieve a relative 3-28% increase in ability to predict downstream performance compared with using scale alone. Analysis of model design decisions reveal insights into data composition, such as the trade-off between language and code tasks at 15-25% code, as well as the negative impact of web data on truthfulness. Broadly, our framework lays a foundation for more systematic investigation of how model development choices shape final capabilities.  [^1] Code and data are available at https://anonymous.4open.science/r/llm-pretraining-behaviours-FE80/
@@
Emmy Liu1, Amanda Bertsch1, Lintang Sutawika1, Lindia Tjuatja1, Patrick Fernandes1,2,3 Lara Marinov1, Michael Chen1, Shreya Singhal1, Carolin Lawrence4, Aditi Raghunathan1, Kiril Gashteovski4, 5, Graham Neubig1 1 Carnegie Mellon University, Language Technologies Institute 2Instituto Superior Técnico (Lisbon ELLIS Unit), 3 Instituto de Telecomunicações 4 NEC Laboratories Europe, Germany 5 CAIR, Ss. Cyril and Methodius University of Skopje, North Macedonia Correspondence: emmy@cmu.edu
@@ 3 Building a Database of Publicly-Available Language Models
To address our research question, we built a database of publicly available language models spanning 11M to 110B parameters, [^3] Including embedding parameters. limited to distinct pretrained decoder-only base models. [^4] By distinct, we mean unique combinations of training data and architecture. Models trained on deduplicated datasets are counted separately, but not variants with different curricula/initializations. This section describes our inclusion criteria, model featurization, and evaluation approach.
@@ 3.2.2 Exploring Data Composition via Generation
Although many models document some data composition details, relatively few release their full pretraining corpus, resulting in missing values for many models in our study.
We propose estimating a model’s training data by generating from it with only a beginning-of-sequence token in context (or end-of-sequence if BOS is unavailable). Using temperature-based sampling with $T=1$,555this should in principle recover $P_{raw}(x_{t}\mid x_{<t})\approx P(x_{t}\mid x_{<t})$ in the limit of sampling infinitely from an LM that captures its distribution perfectly. we sample 5-10k free-generations per model and categorize them using an LM-based classifier (see Appendix E) according to the domains in Figure 2. A human annotator independently validated the classifier on 300 documents from pretraining datasets (Appendix F), showing high agreement (Cohen’s $\kappa=0.746$).
We also extract lower-level linguistic features such as words per sentence, constituency tree depth, and dependency length. Our validation analysis (Appendix G) shows that domain-level features correlate well with actual pretraining data composition (e.g., web content correlation: $r=0.916$, $p=7.55\times 10^{-12}$), while lower-level stylistic features show weaker correlations. However, holistic model-level correlations across all features are strong (typically $r>0.8$), supporting our use of free-generations as proxies for pretraining composition, while not substituting free-generation features for pretraining features.
@@ 3.3 Evaluation Datasets and Metrics
Table 1: Overview of LM evaluation datasets with approximate sample counts and evaluation metrics. Datasets ANLI, XNLI, LogiQA2, and MathQA use Brier Score, while the others use Accuracy.
To assess how design choices affect reasoning capabilities, we evaluated models on datasets from the Open LLM leaderboard Myrzakhan et al. (2024) that capture diverse aspects of reasoning (Table 1). [^6] We excluded Arithmetic and Minerva math tasks Brown et al. (2020); Hendrycks et al. (2021) as we focused on base models, and few achieved non-zero scores. We collect results for some models directly from the leaderboard, and for models not on the leaderboard we use the Eleuther LM eval harness (Gao et al., 2023) to conduct evaluations with exactly the same setting. In addition, if there were multiple versions of a task or sub-tasks, we evaluated all of them and averaged them to get the overall task score. For the full list of evaluation datasets and settings, see Appendix C.
@@
For tasks where smaller models struggle to achieve non-zero accuracy, we follow Schaeffer et al. (2023) in using multiclass brier score as an alternate continuous metric for multiple-choice tasks Brier (1950). [^7] Note that lower is better for brier score. Multiclass brier score ranges between 0-2. For a task with $K$ classes, let $p_{ik}$ be the predicted probability for class $k$ on sample $i$. Then $s_{T,BS}=\frac{1}{|T|}\sum_{i=1}^{|T|}\sum_{k=1}^{K}(p_{ik}-\mathds{1}{\{y_{i}=k\}})^{2}$.
@@ 4 Predictive Modeling
Next, given our database we fit a regressor to try to predict performance. In traditional scaling laws, regressors are fit based on power laws. However, we are now dealing with a larger number of features, some of which may not be captured well by simple parametric forms. Hence, we follow previous work on performance prediction (Xia et al., 2020; Ye et al., 2021) utilizing tree-based regressors based on XGBoost Chen and Guestrin (2016). [^8] We also performed preliminary experiments with LightGBM Ke et al. (2017) but it yielded very similar results in both prediction accuracy and feature importance. The LightGBM version of the main results can be found in Appendix K.
For each evaluation benchmark, we train a model to predict the performance metric on that task based on architectural features $\mathcal{A}$ and data features $\mathcal{D}$. For each task setting, we perform 3-fold cross-validation due to the relatively small number of models, with a nested inner cross-validation over the training set in each fold. The inner cross-validation conducted grid search over a small set of hyperparameters, allowing the model to slightly vary per task. See Appendix I for more details.
@@
To evaluate the predictors, we use Mean Absolute Error averaged across all models and folds. In other words, for a task with $N$ models evaluated, $\text{MAE}_{T}=\frac{1}{|T|}\sum_{i=1}^{N}|s_{T}(M_{i})-\widehat{s_{T}}(M_{i})|$. We compare the scaling-laws predictor as well as the all-features predictor against each other, but also against the median baseline, which simply predicts the median score of the models in the training set for each model in the test set of that fold, and the log-linear baseline, which fits a log-linear function to the number of parameters and number of tokens.
@@
As the full set of features is very large, we sequentially selected features from the full set greedily based on which reduced MAE the most, averaged across 5 random seeds. Features were added until no reduction of at least $1\times 10^{-4}$ was observed. We started using only the two scaling laws features, and refer to this as the scaling-laws model, though it does not have the form of a traditional power law. [^9] As we use a tree-based predictor to accommodate diverse feature types (including non-numeric ones), our approach prioritizes interpolation within observed bounds (10M-100B parameters, 50B-3T tokens) rather than extrapolation. Exploring other prediction methods remains future work. By then incorporating additional architectural or data features, we can then directly quantify the incremental predictive power afforded by these extra features. We refer to the model with the set of features as the all-features model. In all cases, we ran models with the same hyperparameter grid and the same random seeds and splits.
@@
Benchmark Setting Baseline MAE Log-Linear MAE Scaling Laws MAE All Features MAE p-val (corrected)
@@
Arc Challenge 25-shot 13.23% 4.91% 4.36% $\pm$ 0.12% 3.67% $\pm$ 0.09% $3.67\times 10^{-19}$
GSM8k 5-shot 15.65% 11.03% 6.04% $\pm$ 0.21% 5.10% $\pm$ 0.23% $5.41\times 10^{-14}$
Hellaswag 10-shot 12.26% 4.29% 3.93% $\pm$ 0.13% 3.18% $\pm$ 0.09% $4.44\times 10^{-20}$
Humaneval 0-shot 11.79% 8.61% 8.08% $\pm$ 0.22% 6.93% $\pm$ 0.22% $4.44\times 10^{-20}$
Lambada 0-shot 16.89% 9.60% 9.51% $\pm$ 0.33% 6.85% $\pm$ 0.25% $2.87\times 10^{-22}$
MMLU (0-shot) 0-shot 11.98% 9.12% 4.76% $\pm$ 0.20% 4.10% $\pm$ 0.17% $5.26\times 10^{-13}$
MMLU (5-shot) 5-shot 12.25% 8.39% 3.97% $\pm$ 0.18% 3.54% $\pm$ 0.14% $2.09\times 10^{-10}$
TruthfulQA 0-shot 3.72% 3.40% 2.75% $\pm$ 0.08% 2.29% $\pm$ 0.06% $1.02\times 10^{-17}$
Winogrande 5-shot 10.14% 3.99% 3.39% $\pm$ 0.08% 3.09% $\pm$ 0.07% $5.26\times 10^{-13}$
@@
XNLI 0-shot 7.22 5.45 5.11 $\pm$ 0.11% 4.30$\pm$ 0.11% $1.37\times 10^{-2}$
ANLI 0-shot 9.48 5.95 6.18 $\pm$ 0.19% 5.86$\pm$ 0.21% $3.16\times 10^{-9}$
MathQA 0-shot 7.57 3.89 2.83 $\pm$ 0.06% 2.75$\pm$ 0.07% $1.63\times 10^{-4}$
LogiQA2 0-shot 12.62 8.77 4.74 $\pm$ 0.12% 4.60$\pm$ 0.15% $3.84\times 10^{-4}$
Table 2: Comparison of MAE values (mean $\pm$ 95% CI) for Scaling Laws and All Features predictors alongside Baseline MAE and Log-Linear MAE. Lower MAE is bolded; * indicates significance ($p<0.05$). Brier score values are multiplied by 100 to be on a similar scale to accuracy.
@@ 5.2 What Features Does Task Performance Depend On?
To understand factors influencing task performance, we examine Shapley (1953) (SHAP) values, which show how feature values affect predictions. Results for Arc Challenge, HumanEval, Winogrande, and TruthfulQA appear in Figure 4, with remaining benchmarks in Appendix L.
A little code goes a long way, but too much is harmful to NLI. The percentage of code in pretraining is a critical non-scaling feature. Higher code composition benefits Humaneval performance but negatively impacts natural language reasoning tasks including Arc Challenge, Hellaswag, Winogrande, and Lambada. As shown in Figure 5, models with over 20-25% code show gains on Humaneval but penalties on language benchmarks. A moderate 15-25% code proportion appears to balance these competing demands.
@@
Other data domains show task-specific effects. From free-generation features, we observed recent models trained on synthetic data (Phi (Gunasekar et al., 2023), SmolLM (Allal et al., 2024)) generate more question words, suggesting training on question-answering content. Reference-like or question-loaded generations correlate with better performance on Arc Challenge and Winogrande, while web-like generations correlate with worse TruthfulQA performance (Figure 4).
@@
6 Validating Performance Predictions with Confirmatory Experiments
To validate findings from the meta-analysis, we also ran confirmatory pretraining runs with 460M-parameter models on the Dolma dataset. We aimed to validate two data distributional findings: (1) Around 8% code is optimal when only considering natural language inference, but 15-25% may be best when balancing code and natural language, and (2) TruthfulQA performance decreases with an increasing proportion of web data. As this is a small scale model and accuracy differences may not be significant, we convert the relevant datasets to use loss-based evaluations. Due to computational constraints, we train each checkpoint for 10B tokens, but use a cosine learning rate schedule scaled to a 100B token run. See Appendix M for details and exact loss figures. Overall, in Figure 6 find that the confirmatory runs largely validated our meta-analysis predictions, with the exception that our margin-based loss for truthfulQA was slightly lower for the 50% web data checkpoint compared to the 30% checkpoint, though the trend for accuracy is as expected. This provides preliminary evidence that our analysis method could be used to intelligently predict LM training design decisions a-priori.
Figure 6: Loss on 460M parameter models, early checkpoints taken at 10% of final training length. Panel A depicts the effects of code vs. natural language data on natural language datasets and coding, while Panel B depicts the effects of web vs. non-web based data on truthfulQA. In Panel A, the optimal ranges for natural language tasks only (green) and for balancing natural language and code (blue) are highlighted.
@@ 7.1 Empirical Data Composition Results
Prior work has examined code in pretraining Ma et al. (2023); Aryabumi et al. (2024) and domain ablations Longpre et al. (2024). Data filtering improves performance beyond scaling alone Sorscher et al. (2023); Goyal et al. (2024). Our results show code enhances natural language reasoning at moderate proportions (optimal ratio 15-25%), refining previous estimates of 25% Aryabumi et al. (2024). Our approach of pooling insights from existing models complements empirical ablations by identifying promising axes for further testing.
@@ 7.2 Observational and Task-Specific Scaling Law Fitting
Task-specific scaling laws research shows parameter allocation affects machine translation outcomes Ghorbani et al. (2021), and multitasking benefits English-target languages Fernandes et al. (2023). Work on downstream tasks emphasizes alignment between pretraining and downstream data Hernandez et al. (2021); Isik et al. (2024). Various studies address data repetition Muennighoff et al. (2024), multiple domains Goyal et al. (2024), and factors like sparsity Frantar et al. (2023), precision Kumar et al. (2024), and inference costs Hoffmann et al. (2022b), while some find stability across training hyperparameters DeepSeek-AI et al. (2024a).